1. You recently received a birthday gift of $500 and you decide to save it for later use. If you deposit it into an investment account that earns 6 percent interest, how much will be in the account in 3 years?

2. You borrowed $1,200 to buy a new computer and agreed to pay the loan back 3 years later; if the loan has a 10 percent interest rate, what is the balance you will owe at that time?

3. You recently purchased a new car for 18,000 by agreeing to make equal annual payments for 5 years. If the interest rate on your loan is 8 percent, how much is each payment?

4. In 2 years you will need $800 for a trip you planned to Cost Rica. If your account earns 4 percent interest, how much should you deposit today to reach your goal?
Answer #1

1. Assuming Compounding interest: 500*(1+0.06)^3 = 595.51

2. Assuming Compounding annually, Total outstanding at end of year 3: 1200*(1+0.1)^3 = 1597.20

3. Equated annual installments: 18000/(PV For 5 years @ 8%) ie, 18000/3.9927 = 4508.22

4. Assuming Compounding, Amount to be invested: 800=??*(1+0.04)^2 => ?? = 800/(1.04)^2 = 739.65

(Please comment in case of any further doubt regarding the solution.)